OBSERVATION OF THE Ã − X̃ ELECTRONIC TRANSITION FOR VARIOUS PEROXY RADICALS

\underline{SERGEY J. ZALYUBOVSKY}, JEFF GRAY, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus OH 43210.

Cavity ringdown spectra of the near IR Ã − X̃ electronic transition for acetyl, isopropyl and partially fluorinated methyl peroxy radicals are reported. Structured spectral signatures of the Ã − X̃ transition contain information sufficient to distinguish the detailed nature of the R group even when they are quite similar. Therefore cavity ringdown spectroscopy of peroxy (RO_{2}) radicals via the Ã^{2}A′ − X̃^{2}A′′ transition can be successfully used for monitioring various peroxy radicals in complex chemical systems. Ab initio calculations of the origin frequencies for the aforementioned transitions have been carried out at the G2/G2(MP2) level for various peroxy radicals. Remarkable agreement (∼20 cm^{−1}) with observed experimental values of T_{00} for the Ã^{2}A′ − X̃^{2}A′′ transition is observed.